# Long Cheng1 Jiafei Duan1,211footnotemark: 1 Yi Ru Wang1 Haoquan Fang122footnotemark: 2 Boyang Li122footnotemark: 2 Yushan Huang1 Elvis Wang3 Ainaz Eftekhar1,2 Jason Lee1 Wentao Yuan1 Rose Hendrix2 Noah A. Smith1,2 Fei Xia1 Dieter Fox1 Ranjay Krishna1,2

### v1

@@ 1 Introduction
Pointing focuses our attention. It is one of the earliest and most universal non-verbal methods we use to communicate intent; in fact, children learn to point as a prelinguistic form of communication tomasello2007new . Precise spatial grounding—pointing—enables a wide range of practical and high-impact applications across robotics, assistive technology, human-computer interaction, and vision-language interfaces. In robotics, a pointing-capable model can interpret language commands like “pick up the red cup next to the bowl” and translate them into precise spatial actions yuan2025robopoint , enabling fine-grained object manipulation in cluttered environments duan2024manipulate . In assistive technologies, systems can help visually impaired users by answering spatial queries such as “where is the handle on this door?” bigham2010vizwiz or ‘which one is the garlic?’ In education or creative tools, pointing allows for interactive visual tutoring, such as identifying components in scientific diagrams or guiding a learner through a painting hu2024visual . Even in everyday virtual assistants or search engines, the ability to refer to specific image regions via pointing could make multimodal interactions more intuitive and expressive deitke2024molmo . Across these domains, pointing provides a low-bandwidth yet powerful spatial interface for grounding language in vision—precise enough for manipulation ray2024sat , intuitive enough for communication, and general enough to scale with modern multimodal models.
@@
Recent advances in multimodal models have begun to incorporate more dynamic and spatially expressive forms of interaction. The Segment Anything Model (SAM) kirillov2023segment enables segmentation from sparse visual prompts such as points or boxes, revealing the potential of fine-grained spatial control. Google’s Gemini models georgiev2024gemini push the boundaries of long-context visual reasoning, incorporating multiple modalities over extended sequences. In parallel, new datasets have emerged to support explicit spatial referencing. Molmo’s PixMo dataset deitke2024molmo introduces 2D pointing as a form of multimodal alignment between images and instructions, while RoboPoint yuan2025robopoint focuses on spatial affordance prediction by linking instructions to interaction-relevant keypoints in robotic contexts. These setups bias evaluations toward pixel-level accuracy rather than conceptual reasoning and often lack diversity or scalability.
There is a need for a holistic evaluation platform to make progress towards language-guided pointing. Although datasets for referring expressions exist (e.g.RefCOCO, RefCOCO+, and RefCOCOg kazemzadeh2014referitgame ; yu2016modeling ), they are focused on a subset of pointing tasks: object location. They lack the ambiguity and contextual variability that users expect from modern interactive models, limiting their utility for studying pragmatic or interactive applications. As a result, they offer only partial insight into the full spectrum of grounding required for embodied or assistive agents.
@@
PointArena decomposes pointing into three stages of evaluation: 1) Point-Bench is a curated dataset of $982$ manually selected, annotated, and verified image-question pairs across five high-level categories (Spatial, Affordance, Counting, Steerable, and Reasoning), an interactive, online platform for blind, pairwise comparison between models based on user instructions. Users select from curated or custom-uploaded images. Voting is anonymized, and we have collected over $4,500$ votes from more than $100$ participants. 3) Point-Act is a real-world benchmark that evaluates the utility of pointing in for a downstream application. The system directs a robotic arm to manipulate objects through pointing-based language commands. All three evaluation stages require minimal human effort; each is self-contained and can run live to evaluate any model.
Through our evaluation of both open-source and proprietary models across the three stages of the PointArena benchmark, we find that Molmo-72B achieves the highest performance on Point-Bench, with proprietary models such as Gemini-2.5-Pro performing comparably. Models trained with explicit pointing supervision consistently outperform those without. We also observe a strong correlation between static benchmark accuracy and human preference in Point-Battle. Notably, we find that adding language reasoning (e.g., Chain-of-Thought wei2022chain ) does not improve visual grounding for pointing tasks. Our study further reveals several other actionable insights into model behavior and evaluation design. We see PointArena as a missing component necessary as we develop general-purpose vision-language models that can reason about and interact with the world.
@@ 2 Related work
Grounding benchmarks. There are an abundance of benchmarks for visual grounding and spatial reasoning capabilities of multimodal large language models (MLLMs). The RefCOCO, RefCOCO+, and RefCOCOg datasets focus on 2D visual grounding, with RefCOCOg emphasizing long-form referring expressions and fine-grained object distinctions yu2016modelingcontextreferringexpressions . In 3D, ScanRefer provides 51,583 descriptions across 800 RGB-D scans, supporting joint language-geometry models for 3D object localization chen2020scanrefer3dobjectlocalization . ReferIt3D and CityRefer extend this to fine-grained and outdoor settings, with CityRefer incorporating geographic features from OpenStreetMap achlioptas2020referit_3d ; miyanishi2023cityrefergeographyaware3dvisual . Interactive benchmarks such as GuessWhat?! evaluate multi-turn object grounding through binary dialog across 150K games devries2017guesswhatvisualobjectdiscovery . Flickr30K Entities supports phrase-region grounding through 276K bounding boxes and cross-caption coreference annotations plummer2016flickr30kentitiescollectingregiontophrase . These datasets primarily focus on bounding box grounding, object retrieval, or dialog-based localization. They do not explicitly evaluate pointing behavior.
@@
Arena style evaluation. Arena-style evaluations have become a common framework for comparing large language models (LLMs) via pairwise comparisons and user voting. Chatbot Arena established the core methodology, using anonymized head-to-head comparisons and Elo ratings, collecting over 240K human votes across 50+ models and 100+ languages chiang2024chatbotarenaopenplatform . The MT-Bench framework enhanced this approach by introducing a specialized multi-turn dialogue evaluation system that assesses contextual understanding and reasoning through AI-powered scoring zheng2023judgingllmasajudgemtbenchchatbot . am-ELO addressed instability in Elo rankings by incorporating maximum likelihood estimation and annotator reliability modeling liu2025amelostableframeworkarenabased . Auto-Arena fully automated the evaluation process using LLM-generated questions, peer model battles, and LLM-based voting, achieving 92.14% agreement with human votes zhao2024autoarenaautomatingllmevaluations . BenchBuilder introduced automated benchmark construction from crowdsourced data, yielding Arena-Hard-Auto with tighter confidence intervals than MT-Bench. Specialized variants such as Werewolf Arena evaluated models on social reasoning bailis2024werewolfarenacasestudy , while OpenArena enabled offline evaluations using LLMs as judges openarena2024 . Other work has highlighted vulnerabilities, including susceptibility to vote manipulation min2025improvingmodelrankingchatbot , and proposed alternatives like Tournament Evaluation to improve robustness kelley2025tournament . While arena-style methods are scalable and align with user preferences, they lack ground-truth supervision and are susceptible to adversarial influence. PointArena addresses these limitations by combining ground-truth evaluation (Point-Bench) with arena-style human evaluations (Point-Battle).
@@
Models that point or sketch. Multimodal large language models (MLLMs) have shown significant advances in reasoning capabilities, exemplified by GPT-4V, which demonstrates complex vision-language tasks such as image-based story generation and OCR-free mathematical reasoning 10.1093/nsr/nwae403 . Typical MLLM architectures employ encoder-decoder frameworks featuring early, intermediate, or joint visual-language fusion methods to enhance cross-modal integration. For instance, MiniGPT-4 leverages ViT-based visual features aligned with language models such as Vicuna via Q-Formers and learned projection layers wang2024comprehensivereviewmultimodallarge . Molmo notably improves 2D spatial grounding by predicting normalized coordinates directly from modality-specific embeddings, achieving up to 92% precision for tasks like desktop icon localization deitke2024molmopixmoopenweights . RoboPointyuan2024robopoint uses instruction-tuned vision-language models trained on synthetic data to predict robotic affordances, significantly outperforming GPT-4o and PIVOTnasiriany2024pivotiterativevisualprompting by more than 20% in spatial accuracy, and effectively supports tasks in manipulation, navigation, and augmented reality without requiring real-world supervision. Other models, such as VisCPM and Qwen-VL, incorporate region-conditioned controls and multilingual capabilities through multi-stage training approaches 10.1093/nsr/nwae403 . Additionally, newer systems like NExT-GPT extend capabilities further by integrating 3D point clouds, audio, and video modalities 10.1093/nsr/nwae403 . Despite these advancements, significant opportunities remain to further enhance the precision and accuracy of spatial localization within MLLMs. Key open questions include identifying factors that improve pointing capabilities and understanding their underlying mechanisms in multimodal models.
@@ 3 PointArena
Evaluating the ability of MLLMs to localize language-referred entities in images requires benchmarks that are both precise and diagnostic. Existing benchmarks often emphasize classification or captioning, but fall short when it comes to assessing fine-grained spatial grounding—the ability to resolve natural language instructions into specific image coordinates. This capability is critical not only for understanding model alignment with human intent, but also for enabling downstream applications in robotics yuan2024robopoint , augmented reality duan2023ar2 , and interactive web agents gur2023real , and potentially contributing to explainability park2018multimodalexplanationsjustifyingdecisions . As both specialized pointing models and general-purpose MLLMs improve, standardized evaluation across open-source and proprietary systems becomes essential.
@@ 3.2 Point-Bench
Point-Bench is the largest benchmark for evaluating language-guided pointing, offering 982 text-image pairs with pixel-level target masks collected from public sources posted after 20 April 20, 2025. The data set is evenly divided into five task-driven categories, Spatial, Affordance, Counting, Steerable, and Reasoning, identified through a survey of the question types most frequently tackled by open-source MLLMs  deitke2024molmopixmoopenweights ; yuan2024robopoint ; geminiroboticsteam2025geminiroboticsbringingai . Annotators recruited through crowd-sourcing were free to ask any question, but category-specific image selection naturally steers them toward prompts along the desired axis (e.g., scenes with repeated objects for counting). Together, these curated splits enable systematic evaluation of an MLLM’s ability to recognize, reason, and precisely ground language in visual space.
@@
Steerable — Images from the PixMo dataset  deitke2024molmopixmoopenweights each include a reference point. Annotators pose queries relative to that point, avoiding explicit object names; e.g., “Point to the item closest to the marked point.”
@@ 3.3 Point-Battle
As MLLMs increasingly incorporate visually grounded reasoning and pointing capabilities, static benchmarks become inadequate for evaluating performance in open-ended, real-world scenarios—particularly with respect to human preferences. To address this limitation, we introduce Point-Battle, a dynamic platform for pairwise evaluation of MLLMs’ pointing abilities based on user-provided language instructions. Point-Battle adopts a head-to-head evaluation format inspired by Chatbot Arena chiang2024chatbotarenaopenplatform , implemented through a Gradio-based web interface. In each round, two anonymized models are randomly sampled from the top performers in Point-Bench—including GPT-4o, Gemini 2.5 Flash, Molmo-7B-D, Qwen2.5-VL-7B, and Grok-2 Vision. Users submit a natural language instruction and select an image from a curated dataset (post-April 20, 2025) or upload one of their own. The two models return point predictions, which are displayed side-by-side. Participants vote for the better output or select “both good” or “both bad” if applicable. No preset prompts are provided, encouraging diverse and unbiased instructions. Model identities were kept anonymous to prevent bias. Since its launch, Point-Battle has collected over 4,500 votes from approximately 100 participants worldwide. Unlike the static Point-Bench, which may be subject to overfitting if used during model development, Point-Battle serves as a continuously updated benchmark that captures real-time human preferences and tracks progress in visually grounded reasoning across MLLMs. Furthermore, as Point-Battle scales up, this would also be a platform for collecting pointing data.
@@ 4.1 Evaluation Setup
Models. We evaluate variants from Molmo deitke2024molmo , Gemini geminiroboticsteam2025geminiroboticsbringingai , OpenAI achiam2023gpt , Claude anthropic2024claude3 , Grok xai2024grok2 , LLaVA li2024llava , and Qwen bai2025qwen2 . See appendix for details.
@@ 4.2 Main Results
Figure 5: Insights drawn from Point-Battle and Point-Bench. (a) This figure shows Point-Bench performance (%) of MLLMs over time, grouped by model family. A sharp performance increase is observed in models released after the PixMo dataset (dashed line, December 2024). Notably, GPT4.1 improves by 21.1 percentage points over GPT-4-Turbo, and Gemini-2.0-Flash improves by 45.9 points over Gemini-1.5-Flash. These trends suggest that newer proprietary models may incorporate pointing supervision, potentially derived from or inspired by PixMo. (b) Linear regression of the five models common to Point-Battle and Point-Bench reveals a strong correlation ($R^{2}=0.85$), confirming close agreement between the two evaluation frameworks. (c) Performance of open-source models as a function of parameter count. While there is a slight upward trend, the performance gains with increasing model size are marginal, suggesting diminishing returns and limited sensitivity to scale within this range.
Open-source models perform comparably to proprietary models in pointing accuracy. Point-Bench results show that open-source MLLMs explicitly trained on pointing data often match or outperform proprietary models. For example, Molmo-72B outperformed Gemini-2.5-Pro by 0.43 percentage points—a statistically insignificant margin ($p\approx 0.29$). In affordance reasoning, open-source models like Molmo-72B and Qwen2.5-VL consistently exceed proprietary baselines. Overall, Molmo-72B achieves the highest performance on the Point-Bench benchmark as shown in Table 4.
@@
Proprietary models likely benefit from open-source pointing datasets. While proprietary training data is opaque, we observe large performance jumps in models released shortly after the PixMo  deitke2024molmopixmoopenweights and RoboPoint dataset  yuan2024robopoint . For instance, GPT-o3 improved by 21.1 percentage points over GPT-4-Turbo, and Gemini-2.5-Flash improved by 45.9 points over Gemini-1.5-Flash (Figure 5a). These results suggest that recent proprietary models may have incorporated PixMo or a similar corpus.
Open-source models align more closely with human preferences. In Point-Battle, Molmo-7B-D outperformed Gemini-2.5-Flash by 196 Elo points. Their 95% confidence intervals do not overlap, and Molmo-7B-D won 79% of the 115 direct head-to-head comparisons, as shown in Figure 6. Both Qwen2.5-VL-7B and Molmo-7B-D surpass proprietary models in human preference evaluations and exceed the 1000-point baseline, indicating a statistically significant advantage over random guessing. However, in terms of preference-aligned pointing performance, Molmo-7B-D remains clearly superior to Qwen2.5-VL-7B.
@@
Human-preference and static dataset evaluations are highly consistent. Point-Bench’s static dataset will inevitably plateau as MLLMs improve by training on ever-larger, real or synthetic pointing corpora (e.g., RoboPoint yuan2024robopoint ). To stay ahead, we introduce Point-Battle, a live arena that updates continuously and enables open-ended model comparison in real time. Validating this setup, we re-evaluated the models tested on Point-Bench and observed strong alignment: Point-Battle scores correlate with Point-Bench results at $R^{2}=0.85$ (Figure 5b).

### v2

@@ 1 Introduction
Pointing focuses our attention. It is one of the earliest and most universal non-verbal methods we use to communicate intent; in fact, children learn to point as a prelinguistic form of communication Tomasello et al. (2007). Precise spatial grounding—pointing—enables a wide range of practical and high-impact applications across robotics, assistive technology, human-computer interaction, and vision-language interfaces. In robotics, a pointing-capable model can interpret language commands like “pick up the red cup next to the bowl” and translate them into precise spatial actions Yuan et al. (2025), enabling fine-grained object manipulation in cluttered environments Duan et al. (2024). In assistive technologies, systems can help visually impaired users by answering spatial queries such as “where is the handle on this door?” Bigham et al. (2010) or ‘which one is the garlic?’ In education or creative tools, pointing allows for interactive visual tutoring, such as identifying components in scientific diagrams or guiding a learner through a painting Hu et al. (2024). Even in everyday virtual assistants or search engines, the ability to refer to specific image regions via pointing could make multimodal interactions more intuitive and expressive Deitke et al. (2024a). Across these domains, pointing provides a low-bandwidth yet powerful spatial interface for grounding language in vision—precise enough for manipulation Ray et al. (2024), intuitive enough for communication, and general enough to scale with modern multimodal models.
@@
Recent advances in multimodal models have begun to incorporate more dynamic and spatially expressive forms of interaction. The Segment Anything Model (SAM) Kirillov et al. (2023) enables segmentation from sparse visual prompts such as points or boxes, revealing the potential of fine-grained spatial control. Google’s Gemini models Georgiev et al. (2024) push the boundaries of long-context visual reasoning, incorporating multiple modalities over extended sequences. In parallel, new datasets have emerged to support explicit spatial referencing. Molmo’s PixMo dataset Deitke et al. (2024a) introduces 2D pointing as a form of multimodal alignment between images and instructions, while RoboPoint Yuan et al. (2025) focuses on spatial affordance prediction by linking instructions to interaction-relevant keypoints in robotic contexts. These setups bias evaluations toward pixel-level accuracy rather than conceptual reasoning and often lack diversity or scalability.
There is a need for a holistic evaluation platform to make progress towards language-guided pointing. Although datasets for referring expressions exist (e.g.RefCOCO, RefCOCO+, and RefCOCOg Kazemzadeh et al. (2014); Yu et al. (2016a)), they are focused on a subset of pointing tasks: object location. They lack the ambiguity and contextual variability that users expect from modern interactive models, limiting their utility for studying pragmatic or interactive applications. As a result, they offer only partial insight into the full spectrum of grounding required for embodied or assistive agents.
@@
PointArena decomposes pointing into three stages of evaluation: 1) Point-Bench is a curated dataset of $982$ manually selected, annotated, and verified image-question pairs across five high-level categories (Spatial, Affordance, Counting, Steerable, and Reasoning). 2) Point-Battle is an interactive, online platform for blind, pairwise comparison between models based on user instructions. Users select from curated or custom-uploaded images. Voting is anonymized, and we have collected over $4,500$ votes from more than $100$ participants. 3) Point-Act is a real-world benchmark that evaluates the utility of pointing in for a downstream application. The system directs a robotic arm to manipulate objects through pointing-based language commands. All three evaluation stages require minimal human effort; each is self-contained and can run live to evaluate any model.
Through our evaluation of both open-source and proprietary models across the three stages of the PointArena benchmark, we find that Molmo-72B achieves the highest performance on Point-Bench, with proprietary models such as Gemini-2.5-Pro performing comparably. Models trained with explicit pointing supervision consistently outperform those without. We also observe a strong correlation between static benchmark accuracy and human preference in Point-Battle. Notably, we find that adding language reasoning (e.g., Chain-of-Thought Wei et al. (2022)) does not improve visual grounding for pointing tasks. Our study further reveals several other actionable insights into model behavior and evaluation design. We see PointArena as a missing component necessary as we develop general-purpose vision-language models that can reason about and interact with the world.
@@ 2 Related work
Grounding benchmarks. There are an abundance of benchmarks for visual grounding and spatial reasoning capabilities of multimodal large language models (MLLMs). The RefCOCO, RefCOCO+, and RefCOCOg datasets focus on 2D visual grounding, with RefCOCOg emphasizing long-form referring expressions and fine-grained object distinctions Yu et al. (2016b). In 3D, ScanRefer provides 51,583 descriptions across 800 RGB-D scans, supporting joint language-geometry models for 3D object localization Chen et al. (2020). ReferIt3D and CityRefer extend this to fine-grained and outdoor settings, with CityRefer incorporating geographic features from OpenStreetMap Achlioptas et al. (2020); Miyanishi et al. (2023). Interactive benchmarks such as GuessWhat?! evaluate multi-turn object grounding through binary dialog across 150K games de Vries et al. (2017). Flickr30K Entities supports phrase-region grounding through 276K bounding boxes and cross-caption coreference annotations Plummer et al. (2016). These datasets primarily focus on bounding box grounding, object retrieval, or dialog-based localization. They do not explicitly evaluate pointing behavior.
@@
Arena style evaluation. Arena-style evaluations have become a common framework for comparing large language models (LLMs) via pairwise comparisons and user voting. Chatbot Arena established the core methodology, using anonymized head-to-head comparisons and Elo ratings, collecting over 240K human votes across 50+ models and 100+ languages Chiang et al. (2024). The MT-Bench framework enhanced this approach by introducing a specialized multi-turn dialogue evaluation system that assesses contextual understanding and reasoning through AI-powered scoring Zheng et al. (2023). am-ELO addressed instability in Elo rankings by incorporating maximum likelihood estimation and annotator reliability modeling Liu et al. (2025). Auto-Arena fully automated the evaluation process using LLM-generated questions, peer model battles, and LLM-based voting, achieving 92.14% agreement with human votes Zhao et al. (2024). BenchBuilder introduced automated benchmark construction from crowdsourced data, yielding Arena-Hard-Auto with tighter confidence intervals than MT-Bench. Specialized variants such as Werewolf Arena evaluated models on social reasoning Bailis et al. (2024), while OpenArena enabled offline evaluations using LLMs as judges SYV-AI (2024). Other work has highlighted vulnerabilities, including susceptibility to vote manipulation Min et al. (2025), and proposed alternatives like Tournament Evaluation to improve robustness Kelley and Wilson (2025). While arena-style methods are scalable and align with user preferences, they lack ground-truth supervision and are susceptible to adversarial influence. PointArena addresses these limitations by combining ground-truth evaluation (Point-Bench) with arena-style human evaluations (Point-Battle).
@@
Models that point or sketch. Multimodal large language models (MLLMs) have shown significant advances in reasoning capabilities, exemplified by GPT-4V, which demonstrates complex vision-language tasks such as image-based story generation and OCR-free mathematical reasoning Yin et al. (2024). Typical MLLM architectures employ encoder-decoder frameworks featuring early, intermediate, or joint visual-language fusion methods to enhance cross-modal integration. For instance, MiniGPT-4 leverages ViT-based visual features aligned with language models such as Vicuna via Q-Formers and learned projection layers Wang et al. (2024). Molmo notably improves 2D spatial grounding by predicting normalized coordinates directly from modality-specific embeddings, achieving up to 92% precision for tasks like desktop icon localization Deitke et al. (2024b). RoboPointYuan et al. (2024) uses instruction-tuned vision-language models trained on synthetic data to predict robotic affordances, significantly outperforming GPT-4o and PIVOTNasiriany et al. (2024) by more than 20% in spatial accuracy, and effectively supports tasks in manipulation, navigation, and augmented reality without requiring real-world supervision. Other models, such as VisCPM and Qwen-VL, incorporate region-conditioned controls and multilingual capabilities through multi-stage training approaches Yin et al. (2024). Additionally, newer systems like NExT-GPT extend capabilities further by integrating 3D point clouds, audio, and video modalities Yin et al. (2024). Despite these advancements, significant opportunities remain to further enhance the precision and accuracy of spatial localization within MLLMs. Key open questions include identifying factors that improve pointing capabilities and understanding their underlying mechanisms in multimodal models.
@@ 3 PointArena
Evaluating the ability of MLLMs to localize language-referred entities in images requires benchmarks that are both precise and diagnostic. Existing benchmarks often emphasize classification or captioning, but fall short when it comes to assessing fine-grained spatial grounding—the ability to resolve natural language instructions into specific image coordinates. This capability is critical not only for understanding model alignment with human intent, but also for enabling downstream applications in robotics Yuan et al. (2024), augmented reality Duan et al. (2023), and interactive web agents Gur et al. (2023), and potentially contributing to explainability Park et al. (2018). As both specialized pointing models and general-purpose MLLMs improve, standardized evaluation across open-source and proprietary systems becomes essential.
@@ 3.2 Point-Bench
Point-Bench is the largest benchmark for evaluating language-guided pointing, offering 982 text-image pairs with pixel-level target masks collected from public sources posted after 20 April 20, 2025. The data set is evenly divided into five task-driven categories, Spatial, Affordance, Counting, Steerable, and Reasoning, identified through a survey of the question types most frequently tackled by open-source MLLMs  Deitke et al. (2024b); Yuan et al. (2024); Team et al. (2025). Annotators recruited through crowd-sourcing were free to ask any question, but category-specific image selection naturally steers them toward prompts along the desired axis (e.g., scenes with repeated objects for counting). Together, these curated splits enable systematic evaluation of an MLLM’s ability to recognize, reason, and precisely ground language in visual space.
@@
Steerable — Images from the PixMo dataset  Deitke et al. (2024b) each include a reference point. Annotators pose queries relative to that point, avoiding explicit object names; e.g., “Point to the item closest to the marked point.”
@@ 3.3 Point-Battle
As MLLMs increasingly incorporate visually grounded reasoning and pointing capabilities, static benchmarks become inadequate for evaluating performance in open-ended, real-world scenarios—particularly with respect to human preferences. To address this limitation, we introduce Point-Battle, a dynamic platform for pairwise evaluation of MLLMs’ pointing abilities based on user-provided language instructions. Point-Battle adopts a head-to-head evaluation format inspired by Chatbot Arena Chiang et al. (2024), implemented through a Gradio-based web interface. In each round, two anonymized models are randomly sampled from the top performers in Point-Bench—including GPT-4o, Gemini 2.5 Flash, Molmo-7B-D, Qwen2.5-VL-7B, and Grok-2 Vision. Users submit a natural language instruction and select an image from a curated dataset (post-April 20, 2025) or upload one of their own. The two models return point predictions, which are displayed side-by-side. Participants vote for the better output or select “both good” or “both bad” if applicable. No preset prompts are provided, encouraging diverse and unbiased instructions. Model identities were kept anonymous to prevent bias. Since its launch, Point-Battle has collected over 4,500 votes from approximately 100 participants worldwide. Unlike the static Point-Bench, which may be subject to overfitting if used during model development, Point-Battle serves as a continuously updated benchmark that captures real-time human preferences and tracks progress in visually grounded reasoning across MLLMs. Furthermore, as Point-Battle scales up, this would also be a platform for collecting pointing data.
@@ 4.1 Evaluation Setup
Models. We evaluate variants from Molmo Deitke et al. (2024a), Gemini Team et al. (2025), OpenAI Achiam et al. (2023), Claude Anthropic (2024), Grok xAI (2024), LLaVA Li et al. (2024), and Qwen Bai et al. (2025). See appendix for details.
@@ 4.2 Main Results
Figure 5: Insights drawn from Point-Battle and Point-Bench. (a) This figure shows Point-Bench performance (%) of MLLMs over time, grouped by model family. A sharp performance increase is observed in models released after the PixMo dataset or RoboPoint (dashed line, October 2024 or December 2024). Notably, GPT4.1 improves by 21.1 percentage points over GPT-4-Turbo, and Gemini-2.0-Flash improves by 45.9 points over Gemini-1.5-Flash. These trends suggest that newer proprietary models may incorporate pointing supervision, potentially derived from or inspired by RoboPoint or PixMo. (b) Linear regression of the five models common to Point-Battle and Point-Bench reveals a strong correlation ($R^{2}=0.85$), confirming close agreement between the two evaluation frameworks. (c) Performance of open-source models as a function of parameter count. While there is a slight upward trend, the performance gains with increasing model size are marginal, suggesting diminishing returns and limited sensitivity to scale within this range.
Open-source models perform comparably to proprietary models in pointing accuracy. Point-Bench results show that open-source MLLMs explicitly trained on pointing data often match or outperform proprietary models. For example, Molmo-72B outperformed Gemini-2.5-Pro by 0.43 percentage points—a statistically insignificant margin ($p\approx 0.29$). In affordance reasoning, open-source models like Molmo-72B and Qwen2.5-VL consistently exceed proprietary baselines. Overall, Molmo-72B achieves the highest performance on the Point-Bench benchmark as shown in Figure 3.
@@
Proprietary models likely benefit from open-source pointing datasets. While proprietary training data is opaque, we observe large performance jumps in models released shortly after the PixMo  Deitke et al. (2024b) and RoboPoint dataset  Yuan et al. (2024). For instance, GPT-o3 improved by 21.1 percentage points over GPT-4-Turbo, and Gemini-2.5-Flash improved by 45.9 points over Gemini-1.5-Flash (Figure 5a). These results suggest that recent proprietary models may have incorporated PixMo or a similar corpus.
Open-source models align more closely with human preferences. In Point-Battle, Molmo-7B-D outperformed Gemini-2.5-Flash by 196 Elo points. Their 95% confidence intervals do not overlap, and Molmo-7B-D won 79% of the 115 direct head-to-head comparisons, as shown in Figure 6a. Both Qwen2.5-VL-7B and Molmo-7B-D surpass proprietary models in human preference evaluations and exceed the 1000-point baseline, indicating a statistically significant advantage over random guessing. However, in terms of preference-aligned pointing performance, Molmo-7B-D remains clearly superior to Qwen2.5-VL-7B.
@@
Human-preference and static dataset evaluations are highly consistent. Point-Bench’s static dataset will inevitably plateau as MLLMs improve by training on ever-larger, real or synthetic pointing corpora (e.g., RoboPoint Yuan et al. (2024)). To stay ahead, we introduce Point-Battle, a live arena that updates continuously and enables open-ended model comparison in real time. Validating this setup, we re-evaluated the models tested on Point-Bench and observed strong alignment: Point-Battle scores correlate with Point-Bench results at $R^{2}=0.85$ (Figure 5b).
@@
Method Affordance Spatial Reasoning Steerability Counting Average
GPT-4o + In-Context (2-shots) 46.0 26.7 23.9 22.5 33.2 30.4
GPT-4o + Chain-of-Thought (CoT) Wei et al. (2022) 41.4 24.6 21.2 13.5 32.1 26.6
GPT-4o + Unparsed language instruction 37.8 23.7 19.7 21.9 31.1 26.9
GPT-4o (Default) 42.4 25.6 23.8 24.5 31.1 29.5
Gemini-2.5-Flash + Chain-of-Thought (CoT) Wei et al. (2022) 44.4 25.6 24.9 26.0 33.7 30.9
Gemini-2.5-Flash (Default) 58.6 47.7 43.2 35.0 49.7 46.8